Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Deck of cards

Expert replies
by Nailya » Mon Mar 30, 2009 3:40 pm
Bill has a small deck of 12 playing cards made up of only 2 suits of 6 cards each. Each of the 6 cards within a suit has a different value from 1 to 6; thus, there are 2 cards in the deck that have the same value.

Bill likes to play a game in which he shuffles the deck, turns over 4 cards, and looks for pairs of cards that have the same value. What is the chance that Bill finds at least one pair of cards that have the same value?

[spoiler]OA 17/33[/spoiler]

Please explain how to solve this problem quickly?
GMATPrep1 - 550
GMAT Prep2 - 590
MGMAT1 - 620
MGMAT2 - 640
MGMAT3 - 660
MGMAT4 - 680
MGMAT5 - 690
GMATPrep1 - 690
MGMAT6 - 720
Join the discussion
Source: — Problem Solving |

by mike22629 » Mon Mar 30, 2009 5:50 pm
The best way to approach this problem is to find the probability of finding no pairs and subtracting that from one.

Hence:
Probability of picking first: 1 (because it can be any card)
Probability of second card not pairing: 10/11
Probability of third card not pairing : 8/10
Probability of fourth card not pairing: 6/9

1*(10/11)(8/10)(6/9) = 16/33 (probability of no pairs)

1 - (16/33) = 17/33

Probability of at least one pair is 17/33
Join the discussion

Re: Deck of cards

by Vemuri » Mon Mar 30, 2009 6:14 pm
Probability of selecting first card = 12/12 = 1 ( It can be any card).

Probability of selecting second card = 10/11 ( 11 cards are left for selection. We shouldn't select the matching card. )

Probability of selecting the third card = 8/10 ( 10 cards left we shouldn't select the two cards that match the already selected 2 cards)

Probability of selecting the 4th card = 6/9 ( 9 cards are left, we shouldn't select the 3 cards that match the already selected 3 cards)

Probability of selecting the cards with no pairs = 1* 10/11*8/10*6/9 = 16/33

Probability of selecting at least one pair = 1-16/33 = 17/33.
Join the discussion

by Nailya » Wed Apr 01, 2009 9:46 am
Thank you guys!!! :D
Join the discussion